# Turning the Tide on Dark Pools? Towards Multi-Stakeholder Vulnerability Notifications in the Ad-Tech Supply Chain

### v1

@@ 2.3 Vulnerability notification campaigns
Notifications of security and privacy vulnerabilities. Researchers have studied the effectiveness of notifications for warning server operators about unintended abuse of their infrastructure for malicious purposes [27, 36]. They have also examined how well these notifications work for fixing various security vulnerabilities, such as HTTPS misconfigurations [37], issues in version control system (VCS) repositories [30, 38], firewall omissions for IPv6 services [25], DDoS amplification vulnerabilities [39, 25], Heartbleed [40], cross-site scripting (XSS) bugs [24], and others [41, 31, 42, 29, 28, 26, 43]. More recent, related research has studied the remedial effects of privacy non-compliance notifications [44, 45, 46, 41]. Despite extensive research on notification campaigns, most studies focus on single-stakeholder scenarios where one entity can resolve a vulnerability. Ours is the first to evaluate notification campaigns in the ad-tech supply chain and any multi-stakeholder scenario.
@@
Notification campaigns in the online advertising supply chain. Although previous research has extensively studied ad-tech supply chain vulnerabilities, it has not examined the effectiveness of large-scale notifications in remedying these vulnerabilities. The work of Vekaria et al. [8] is closest to this study; they conducted a small-scale disclosure of dark pooling vulnerabilities to 55 reputable brands, treating it as a single-stakeholder issue and ignoring the multi-stakeholder aspect of the supply chain. In contrast, we build an effective and large-scale multi-stakeholder vulnerability notification framework for the online advertising supply chain. By focusing on dark pooling, we study the influence and attitudes of each stakeholder (publishers, ad-networks, and advertisers) towards notifications sent from academic researchers and a well-known ad-tech watchdog group, CheckMyAds [34] (i.e., activists). Our automated framework for detecting vulnerabilities and sending notifications is publicly available at $<$URL blinded for review$>$, as is the (non-email) data that is the basis of this paper.
@@
Challenges in selecting notification delivery mechanisms. Notification studies often involve the large-scale detection and notification of vulnerabilities [47, 48, 49]. A common challenge is reliably delivering notifications to the intended recipients. Often, notifications go undelivered due to the lack of publicly known communication channels. Previous research has explored various non-automated notification channels, such as telephone, contact forms on websites, social media accounts, physical mail, and manually identified contact email addresses [38, 31, 30]. These manual approaches achieve high delivery success but require significant manual effort, making them infeasible for large-scale studies like ours, which focuses on the globally distributed multi-stakeholder ad-tech supply chain To address this challenge, researchers have used various automated alternatives for delivering notifications, such as sending them to generic email addresses, or obtaining addresses from hosting providers [27], WHOIS contact information [37, 27, 25, 28, 38, 40, 24, 50], trusted third parties like CERTs or clearing houses [24, 25, 28, 39], and DNS operators [28]. However, these email-based automated approaches often suffer from low delivery success rates due to outdated contact details [24, 28, 27], spam filters [24, 38], or incorrect recipients who do not forward the email to the responsible entity [25]. Additional challenges such as recipients not trusting the senders also emerge [42, 29, 38, 37]. In general, prior research shows a trade-off between scale and deliverability of notifications.
@@ 3 Research Methods
In this section, we describe our methods for identifying entities involved in dark pooling Section 3.1 and the multi-stakeholder vulnerability notification campaign design Section 3.2.
@@ 3.1 Identifying dark pooling vulnerabilities
Curating problematic publishers. Problematic publishers are more than twice as likely to be pooled compared to other publishers [8]. Therefore, we focus on this subset of publishers. Table I lists the seven categories of problematic publishers in our dataset. We started by collecting a list of publishers already classified as problematic by prior research. Since identifying these publishers is not our main goal, we used these existing classifications. First, we removed duplicates and checked if the publishers’ websites were functional. We discarded 79,938 publishers that were either non-functional or had parked domains. Next, we used an advertising filter list [51] to see if the remaining 12,875 publishers had advertising. As expected, many problematic publishers (like typosquatted publishers) did not have ads. We found at least one advertising request on 1,125 of these publishers. To avoid including legitimate publishers and to confirm the presence of ads, we manually checked the content on these publishers. We kept 684 problematic publishers that had one or more display ads on their homepage. We also included one random subpage from these 684 publishers and from the remaining 441 publishers if the subpage served display ads. Our final dataset contains 1,478 URLs across 1,007 distinct publishers.
@@
Misinformation/ [8] (669); [52] (3K); 660
Disinformation [53] (11); [54] (1.6K); [55] (2.1K)
Typosquatting [50] (10.5K) 14
Phishing [50] (72.1K) 309
Piracy [56] (2.7K) 136
Sanctioned [57] (53) 6
@@
Discovering vulnerabilities with static analysis. Here, we use static standards files (i.e., ads.txt and sellers.json) to discover dark pooling vulnerabilities.
@@
Curating a dataset of standards files. First, we crawl ads.txt files from the root of each problematic publishers we are studying, as well as from the Tranco Top-1M websites [58]. We extract the distinct domains of ad-networks listed in these ads.txt files (i.e., domains associated with ads.txt DIRECT and RESELLER entries). Next, we crawl the corresponding sellers.json files from the root of each previously identified seller domain. We extract the distinct ad-network domains listed in these sellers.json files (i.e., domains associated with sellers.json INTERMEDIARY and BOTH entries). We repeat this process until no new seller domains/entities are discovered. This ensures complete coverage of all supply chain entities involved in the sale of ad inventory on our problematic publishers and the Tranco top websites.
@@
Identifying dark pools from the standards files. We use the above data to identify dark pools by finding all publishers whose ads.txt files share a common seller ID with another publisher on some ad-network. We use the sellers.json file of the corresponding ad-network to determine the owner of such pooled seller IDs (i.e., the owner domain). The ad-network that allows such pooling is referred to as the pooled domain. To determine if a pool of publishers sharing a seller ID is a dark pool, we use DuckDuckGo’s entity list [59] to map each publisher in the pool to its parent organization. If a pool contains publishers owned by more than one parent organization and includes at least one of the problematic publishers listed in Table I, then it is classified as dark pool. In total, our static analysis yielded 26.1K dark pools. These pools involved 399 problematic publishers, 1.7K unique victim publishers, and 962 unique ad-networks.
@@
Discovering vulnerabilities with dynamic analysis. As highlighted in prior work [8], static analysis alone cannot prove that dark pooling is actually occurring. This is because each publisher is responsible only for the content of their own ads.txt files and have no control over the misrepresentations in other publishers’ ads.txt files. Therefore, we also perform dynamic analysis to gather evidence that a problematic publisher is actually monetizing its ad inventory using another publisher’s seller ID.
@@
Identifying entities associated with dark pools from live page loads. We follow the methodology developed by Vekaria et al. [8]. Following the crawling configuration and disclosures recommendations of Ahmad et al. [60], we used a stateless web crawler driven by Selenium (v4.1.0) and the Chrome browser (v117.0) with bot mitigation strategies and Xvfb from a non-cloud vantage point to crawl problematic publishers and capture HTTP archive (HAR) files. During each dynamic crawl, we load the problematic publisher, we wait 30 seconds for all resources, including ads, to finish loading. We click on the DOM elements associated with each ad URL and wait for the advertiser’s website to be loaded. We log the associated URL, the chain of redirects, requests, responses, and payloads using the HAR file format. Then, from our HAR files, we extract any strings which have a ‘key:value’ or ‘key=value’ format. We examine if any of the seller IDs associated with our static dark pools appear in these extracted pairs. If they do, we log the crawled website (publisher), the pooled seller ID, the ad-network that issued the pooled seller ID, the domain to which the pooled seller ID was issued to (owner domain), and the advertiser whose creative was served. In a normal transaction, we expect that the owner domain matches the crawled publisher website, or that they are at least organizationally-related. However, in our current scenario examining dark pools containing problematic publishers, the owner domains are (victim) publishers who are not associated with the problematic publishers. In total, our dynamic analysis yielded 1.3K dark pools which included 200 unique ad-networks, 347 unique owner domains, and 889 unique pooled seller IDs. The problematic publishers in these dark pools were observed with ads from 635 unique advertisers.
@@ 3.2 Designing multi-stakeholder notifications
Identifying notification recipients. During our study, we sent notifications to the three types of stakeholders (cf. Section 2.1) in three separate rounds. This allowed us to measure each stakeholder group’s influence in remedying the dark pooling vulnerability accurately.
Round 1: Notifying victim publishers. We notified all popular (i.e., Tranco top-10K) publishers involved in dark pools with one of our problematic publishers, identified through our static analysis. We identified 1.7K such publishers.
@@
Ordering of stakeholders. This is a crucial design choice that multi-stakeholder notifications should evaluate as it can impact how different stakeholders respond and remediate vulnerabilities. In our case, the ordering of notifications was decided based on our intuition about which stakeholders would be most influential at remedying vulnerabilities. We hypothesize that notifications to victim publishers are least influential because they are not monetarily impacted (at least directly) by the vulnerability and also do not have the ability to influence pool memberships directly. Ad-networks, on the other hand, have the ability to re-assign pool memberships when they are the issuers of the pooled seller ID. However, doing so has the potential to lower the revenue generated by their pooled inventory. In contrast, we hypothesize that notifications to advertisers will be most influential since they are the most affected stakeholder, from monetary and reputation perspectives. Additionally, their ability to control the flow of revenue to other stakeholders also makes them more capable of influencing their actions. Therefore, our ordering (notifying publishers, then ad-networks, and then advertisers) is expected to leave enough unresolved dark pools to allow for valid statistical analysis in each round.
@@
Source 3: sellers.json files. We extracted ad-network email addresses from their sellers.json files. These addresses are extracted from the contact__\__email key in the sellers.json file. Per the sellers.json standard, this email address may be used to contact the Advertising System for questions or inquiries about the associated sellers.json file.
Source 4: Common email prefixes. As a final resort, we also consider email prefixes commonly seen in contact emails by different companies. These included info@, support@, help@, webmaster@, and contact@. If the other methods did not provide an email address, we tested the availability of these common addresses. To ensure deliverability of our notifications, we sent test emails from an alternate account and monitored for bounce messages, preserving the sender reputation of the main email account used for the notifications.
@@
Treatment 2: Activists as notification sources ($\mathcal{T}_{2}$). We collaborated with the Check My Ads (CMA) institute [34] (previously known as “Sleeping Giants” [61, 33]), a well-known activist organization which names-and-shames ad-tech entities that aid the monetization of problematic content. Unfortunately, due to organizational limitations (cf. Section 6.2), we couldn’t use CMA’s servers for sending these emails and used our university servers instead. However, recipients in this group received notifications which had CMA branding and explicitly mentioned CMA’s involvement.
@@
The project webpage. The webpage contained a description of our research project’s goals and dark pool identification methodology. This page also contained FAQ and Contact sections to help recipients better understand their notifications or reach us. In our descriptions and notifications, we consciously avoid alarmist or accusatory language. For example, we use the term ‘ad safety research’ in lieu of ‘ad fraud research’ and consistently describe the identified vulnerabilities as ‘potential vulnerabilities’ and problematic publishers as ‘potentially unsafe websites’. This is done to develop a positive relationship with the entities and show intent to aid vulnerability resolution.
@@
Body of notification emails. In addition to an introductory preamble described earlier, the body of notification emails contained summaries of the vulnerabilities identified by our research. We crafted unique summaries for each stakeholder.
  • Victim publishers. We report vulnerabilities (i.e., from static and dynamic analysis) using the following text summaries: (1) For vulnerabilities detected via dynamic analysis: “During network traffic analysis of problematic publishers, we observed that <Number> sellerID(s) in your ads.txt are being used by <Number> potentially problematic publisher(s) to monetize their ad inventory.”; and (2) For vulnerabilties detected via static analysis: “We observed that <Number> sellerID(s) in your ads.txt are being used in ads.txt of at least <Number> other potentially problematic publishers.”
  • Ad-networks. We report two types of vulnerabilities from each of our analysis approaches using the following text summaries: (1) For vulnerabilities detected by static analysis of ads.txt files: “<Number> seller IDs issued by you are being pooled by <Number> potentially problematic publishers.”; (2) For vulnerabilities detected by static analysis of sellers.json files: “<Number> seller IDs owned by you and issued by another ad-network are being pooled by <Number> potentially problematic publishers.”; (3) For ads.txt vulnerabilities confirmed by dynamic analysis: “We confirmed that <Number> seller IDs issued by you are being pooled by <Number> potentially problematic publishers.”; and (4) For sellers.json vulnerabilities confirmed by dynamic analysis: “We confirmed that <Number> seller IDs owned by you and issued by another ad-network are being pooled by <Number> potentially problematic publishers”.
  • Advertisers. Because we do not expect advertisers to be particularly technically adept, we frame our notification around the presence of their ad creatives on potentially problematic content as follows: “An ad creative associated with your brand was observed on <Number> problematic publishers. This association could negatively impact your reputation and future business.”
Reports attached to notification emails. To provide concrete evidence to support the reported vulnerabilities, we generated an automated PDF report, with appropriate branding based on the notification source. This report contained: (1) up to 25 instances of vulnerabilities which involve the specific entity; (2) an explanation of the root causes and implications of each vulnerability; and (3) possible remediation options. In addition, notifications to advertisers also included attachments of screenshots which showed the advertiser’s ad creative on a problematic publisher and the HAR file associated with the page load.
@@
Timeline of notifications. Figure 3 depicts the timeline associated with our notification campaign. Prior to each round of notifications we conduct static and dynamic analysis to identify dark pooling vulnerabilities which involve our dataset of problematic publishers (cf. Section 3.1). Then, we generate and send our vulnerability notifications and reports for each entity associated with that round. If we do not hear back, we send a reminder email approximately 1-2 weeks after the initial email. We then wait a month before once more conducting our static and dynamic analysis. We use the data from this analysis to assess the influence of notifications sent during the current round.

### v2

@@ 2.3 Vulnerability notification campaigns
Notifications of security and privacy vulnerabilities. Researchers have studied the effectiveness of notifications for warning server operators about unintended abuse of their infrastructure for malicious purposes [cetin2016understanding, vasek2012malware]. They have also examined how well these notifications work for fixing various security vulnerabilities, such as HTTPS misconfigurations [zeng2019fixing], issues in version control system (VCS) repositories [maass2021snail, stock2018didn], firewall omissions for IPv6 services [li2016you], DDoS amplification vulnerabilities [kuhrer2014exit, li2016you], Heartbleed [durumeric2014matter], cross-site scripting (XSS) bugs [stock2016hey], and others [maass2021best, maass2021effective, ccetin2018let, ccetin2019tell, cetin2017make, li2016remedying, canali2013role]. More recent, related research has studied the remedial effects of privacy non-compliance notifications [stover2023website, soussi2020feasibility, utz2023comparing, maass2021best]. Despite extensive research on notification campaigns, most studies either focus on single-stakeholder scenarios or multi-stakeholder ones where different stakeholders are notified about a vulnerability that exist at a single stakeholder [cetin2017make] or can be fixed by a single stakeholder [stock2016hey] or that has clear aligned incentives for all stakeholders [li2016you]. Ours is the first to evaluate notification campaigns in the fragmented multi-stakeholder ad-tech supply chain where incentives, responsibilities, vulnerabilities and fixes need cooperation and effort from all stakeholders.
@@
Notification campaigns in the online advertising supply chain. Although previous research has extensively studied ad-tech supply chain vulnerabilities, it has not examined the effectiveness of large-scale notifications in remedying these vulnerabilities. The work of Vekaria et al. [vekaria2023inventory] is closest to this study; they conducted a small-scale disclosure of dark pooling vulnerabilities to 55 reputable brands, treating it as a single-stakeholder issue and ignoring the fragmented multi-stakeholder aspect of the ad-tech supply chain. In contrast, we build an effective and large-scale multi-stakeholder vulnerability notification framework for the online advertising supply chain. By focusing on dark pooling, we study the influence and attitudes of each stakeholder (publishers, ad-networks, and advertisers) towards notifications sent from academic researchers and a well-known ad-tech watchdog group, CheckMyAds [checkmyads] (i.e., activists). Our automated framework for detecting vulnerabilities and sending notifications is publicly available at <<URL blinded for review>>, as is the (non-email) data that is the basis of this paper.
@@
Challenges with notification delivery mechanisms. Notification studies often involve the large-scale detection and notification of vulnerabilities [durumeric2013zmap, cui2011reflections, bajpai2019dagstuhl]. A common challenge is reliably delivering notifications to the intended recipients. Often, notifications go undelivered due to the lack of publicly known communication channels. Previous research has explored various non-automated notification channels, such as telephone, contact forms on websites, social media accounts, physical mail, and manually identified contact email addresses [stock2018didn, maass2021effective, maass2021snail]. These manual approaches achieve high delivery success but require significant manual effort, making them infeasible for large-scale studies like ours, which focuses on the globally distributed multi-stakeholder ad-tech supply chain. To address this challenge, researchers have used automated alternatives for notifications, such as targeting generic email addresses, or obtaining addresses from hosting providers [cetin2016understanding], WHOIS contact information [zeng2019fixing, cetin2016understanding, li2016you, cetin2017make, stock2018didn, durumeric2014matter, stock2016hey, pouryousef2020extortion], trusted third parties like CERTs or clearing houses [stock2016hey, li2016you, cetin2017make, kuhrer2014exit], and DNS operators [cetin2017make]. However, these email-based automated approaches often suffer from low delivery success rates due to outdated contact details [stock2016hey, cetin2017make, cetin2016understanding], spam filters [stock2016hey, stock2018didn], or incorrect recipients who do not forward the email to the responsible entity [li2016you]. Additional challenges such as lack of trust in senders also emerge [ccetin2018let, ccetin2019tell, stock2018didn, zeng2019fixing]. In general, prior works show a trade-off between scale and notification deliverability.
@@ 3.1 Identifying dark pooling vulnerabilities
Curating problematic publishers. Problematic publishers are more than twice as likely to be pooled compared to other publishers [vekaria2023inventory]. Therefore, we focus on this subset of publishers. Table I lists the seven categories of problematic publishers in our dataset. We started by collecting a list of publishers already classified as problematic by prior research. Since identifying these publishers is not our main goal, we used these existing classifications. First, we removed duplicates and checked if the publishers’ websites were functional. We discarded 79,938 publishers that were either non-functional or had parked domains. Next, we used an advertising filter list [easylist] to see if the remaining 12,875 publishers had advertising. As expected, many problematic publishers (like typosquatted publishers) did not have ads. We found at least one advertising request on 1,125 of these publishers. To avoid including legitimate publishers and confirm the presence of ads, we manually checked the publisher content. We kept 684 problematic publishers that had one or more display ads on their homepage. We also included one random subpage from these 684 publishers and from the remaining 441 publishers if the subpage served display ads. Our final dataset contains 1,478 URLs across 1,007 publishers.
@@
Misinformation/ [vekaria2023inventory] (669); [misinfo-dataset] (3K); 660
Disinformation [climate-disinfo-dataset] (11); [moore2023consumption] (1.6K); [fakenews-dataset] (2.1K)
Typosquatting [pouryousef2020extortion] (10.5K) 14
Phishing [pouryousef2020extortion] (72.1K) 309
Piracy [piracy-dataset] (2.7K) 136
Sanctioned [ofac-list] (53) 6
@@
Discovering vulnerabilities with static analysis. We use static files of the standards (i.e., ads.txt and sellers.json) to detect dark pooling vulnerabilities.
@@
Curating a dataset of standards files. First, we crawl ads.txt files from the root of each problematic publishers we are studying, as well as from the Tranco Top-1M websites [pochat2018tranco]. We extract the distinct domains of ad-networks listed in these ads.txt files (i.e., domains associated with ads.txt DIRECT and RESELLER entries). Next, we crawl the corresponding sellers.json files from the root of each previously identified seller domain. We extract the distinct ad-network domains listed in these sellers.json files (i.e., domains associated with sellers.json INTERMEDIARY and BOTH entries). We repeat this process until no new seller domains/entities are discovered. This ensures complete coverage of all supply chain entities involved in the sale of ad inventory on our problematic publishers and the Tranco top websites.
@@
Identifying dark pools from the standards files. We use the above data to identify dark pools by finding all publishers whose ads.txt files share a common seller ID with another publisher on some ad-network. We use the sellers.json file of the corresponding ad-network to determine the owner of such pooled seller IDs (i.e., the owner domain). The ad-network that allows such pooling is referred to as the pooled domain. To determine if a pool of publishers sharing a seller ID is a dark pool, we use DuckDuckGo’s entity list [DDG-trackerradar] to map each publisher in the pool to its parent organization. To ensure 100% completeness and accuracy, we manually validated organizational relationships of the small number entities that weren’t auto-matched in the DuckDuckGo’s list. If a pool contains publishers owned by more than one parent organization and includes at least one of the problematic publishers listed in Table I, then it is classified as dark pool. In total, our static analysis yielded 26.1K dark pools. These pools involved 399 problematic publishers, 1.7K unique victim publishers, and 962 unique ad-networks.
@@
Discovering vulnerabilities with dynamic analysis. As highlighted in prior work [vekaria2023inventory], static analysis alone cannot prove that dark pooling is actually occurring. This is because each publisher is responsible only for the content of their own ads.txt files and have no control over the misrepresentations in other publishers’ ads.txt files. Therefore, we also perform dynamic analysis to gather evidence that a problematic publisher is actually monetizing its ad inventory using another publisher’s seller ID.
@@
Identifying entities associated with dark pools from live page loads. We follow the methodology developed by Vekaria et al. [vekaria2023inventory]. Following the crawling configuration and disclosures recommendations of Ahmad et al. [ahmad2020apophanies], we used a stateless web crawler driven by Selenium (v4.1.0) and the Chrome browser (v117.0) with bot mitigation strategies and Xvfb from a non-cloud vantage point to crawl problematic publishers and capture HTTP archive (HAR) files. During each dynamic crawl, we load the problematic publisher, we wait 30 seconds for all resources, including ads, to finish loading. We click on the DOM elements associated with each ad URL and wait for the advertiser’s website to be loaded. We log the associated URL, the chain of redirects, requests, responses, and payloads using the HAR file format. Then, from our HAR files, we extract any strings which have a ‘key:value’ or ‘key=value’ format. We examine if any seller ID associated with our static dark pools appear in these extracted pairs. If they do, we log the crawled website (publisher), the pooled seller ID, the ad-network that issued the pooled seller ID, the domain to which the pooled seller ID was issued to (owner domain), and the advertiser whose ad was served. In a normal transaction, we expect that the owner domain matches the crawled publisher website, or that they are at least organizationally-related. However, in our current scenario examining dark pools containing problematic publishers, the owner domains are (victim) publishers who are not associated with the problematic publishers. In total, our dynamic analysis yielded 1.3K dark pools which included 200 unique ad-networks, 347 unique owner domains, and 889 unique pooled seller IDs. The problematic publishers in these dark pools were observed with ads from 635 unique advertisers.
Definition of vulnerability and its remediation. In our study, a “vulnerability” is a potential for unauthorized use or abuse of advertising ID(s) by malicious entities, and a vulnerability is considered “fixed” or “remediated” if we detect that the vulnerable advertising ID of a given entity is no longer abused by malicious entities. We consider any form of (partial or complete) reduction in vulnerabilities as remediation. This is because complete removal may not always be guaranteed due to multiple external factors such as business contracts in-place or internal decisions. These are beyond our scope/control and we acknowledge as a limitation (Section 6.2)
@@ 3.2 Designing multi-stakeholder notifications
Identifying notification recipients. We sent notifications to the three types of stakeholders (cf. Section 2.1) in three separate rounds. This allowed us to accurately measure each stakeholder group’s influence in remedying the vulnerabilities.
Round 1: Notifying victim publishers. We notified all popular (Tranco top-10K) publishers involved in dark pools with one of our problematic publishers, identified through our static analysis. We identified 1.7K such publishers.
@@
Ordering of stakeholders. This is a crucial design choice that multi-stakeholder notifications should evaluate as it can impact how different stakeholders respond and remediate vulnerabilities. In our case, the ordering of notifications was decided based on our intuition about which stakeholders would be the most influential at remedying vulnerabilities. We hypothesize that notifications to victim publishers are least influential because they are not monetarily impacted (at least directly) by the vulnerability and also do not have the ability to influence pool memberships directly. Ad-networks, on the other hand, have the ability to re-assign pool memberships when they are the issuers of the pooled seller ID. However, doing so has the potential to lower the revenue generated by their pooled inventory. In contrast, we hypothesize that notifications to advertisers will be the most influential since they are the most affected stakeholder, from monetary and reputational perspectives. Ahmad et al. [ahmad2024companies] shows that brand-safety most directly impacts advertisers when their ad ends up on a problematic website. Advertisers face direct consumer backlash in the form of boycotts and negative press attention as a result of which they are incentivized to not advertise on such publishers, corroborating our hypothesis. Additionally, their ability to control the flow of revenue to other stakeholders also makes them more capable of influencing their actions. Therefore, our ordering (notifying publishers, then ad-networks, and finally advertisers) is expected to leave enough unresolved dark pools to allow for validity and statistical rigor during analysis in each round.
Design choice of parallel vs. sequential notifications We adopt sequential multi-round approach for notifications as opposed to simultaneous or parallel notification approach due to a strong interconnectedness between different stakeholders. Typically publishers are listed on “multiple” popular ad-exchanges and advertisers typically buy ads across “multiple” popular ad-exchanges – making it impossible to attribute the remediation to a specific notification, if notified simultaneously. For example, notifications to a subset of publishers could result in them reaching out to the respective ad-exchanges they work, who may also have been separately notified in the parallel notification strategy. In this case, it would be impossible to attribute remediation to a specific publisher or ad-exchange. We were unable to partition publishers, ad-exchanges, and advertisers into disjoint control and treatment sets so as to avoid such “contamination” in the parallel notification approach. However, in the sequential multi-round approach, our sequential notifications across different rounds focus on notifying unresolved vulnerabilities in each round, which allows us to easily attribute a remediation to the stakeholder in the latter round. The time gap of several weeks between different rounds mitigates the risk of contamination. There remains a non-zero risk but it is much smaller in sequential notifications than in parallel notifications. We further discuss this as part of the limitations (cf. Section 6.2).
@@
Source 3: sellers.json files. We extracted ad-network email addresses from their sellers.json files. These addresses are extracted from the contact_\_email key in the sellers.json file. Per the sellers.json standard, this email address may be used to contact the Advertising System for questions or inquiries about the associated sellers.json file.
Source 4: Common email prefixes. As a final resort, we also consider email prefixes commonly seen in contact emails by different companies. These included info@, support@, help@, webmaster@, and contact@. If the other methods did not provide an email address, we tested the availability of these common addresses. To ensure deliverability of our notifications, we sent test emails from an alternate account to monitor for bounces, preserving sender reputation of the main account used for the notifications.
@@
Treatment 2: Activists as notification sources ($\mathcal{T}_{2}$). We collaborated with the Check My Ads (CMA) institute [checkmyads] (previously known as “Sleeping Giants” [li2021beyond, braun2019activism]), a well-known activist organization which names-and-shames ad-tech entities that aid the monetization of problematic content. Unfortunately, due to organizational limitations (cf. Section 6.2), we couldn’t use CMA’s servers for sending these emails and used our university servers instead. However, recipients in this group received notifications which had CMA branding and explicitly mentioned CMA’s involvement.
@@
The project webpage. The webpage (ad-inventory-fraud.github.io) contained a description of our research project’s goals and dark pool identification methodology. This page also contained FAQ and Contact sections to help recipients better understand their notifications or reach us. In our descriptions and notifications, we consciously avoid alarmist or accusatory language. For example, we use the term ‘ad safety research’ in lieu of ‘ad fraud research’ and consistently describe the identified vulnerabilities as ‘potential vulnerabilities’ and problematic publishers as ‘potentially unsafe websites’. This is done to develop a positive relationship with the entities and show intent to aid vulnerability resolution.
@@
Body of notification emails. In addition to an introductory preamble described earlier, the email body contained summaries of the vulnerabilities identified by our research. We crafted unique summaries for each stakeholder.
Victim publishers. We report vulnerabilities (i.e., from static and dynamic analysis) using the following text summaries: (1) For vulnerabilities detected via dynamic analysis: “During network traffic analysis of problematic publishers, we observed that <num> sellerID(s) in your ads.txt are being used by <num> potentially problematic publisher(s) to monetize their ad inventory.”; and (2) For vulnerabilties detected via static analysis: “We observed that <num> sellerID(s) in your ads.txt are being used in ads.txt of at least <num> other potentially problematic publishers.”
Ad-networks. We report two types of vulnerabilities from each of our analysis approaches using the following text summaries: (1) For vulnerabilities detected by static analysis of ads.txt files: “<num> seller IDs issued by you are being pooled by <num> potentially problematic publishers.”; (2) For vulnerabilities detected by static analysis of sellers.json files: “<num> seller IDs owned by you and issued by another ad-network are being pooled by <num> potentially problematic publishers.”; (3) For ads.txt vulnerabilities confirmed by dynamic analysis: “We confirmed that <num> seller IDs issued by you are being pooled by <num> potentially problematic publishers.”; and (4) For sellers.json vulnerabilities confirmed by dynamic analysis: “We confirmed that <num> seller IDs owned by you and issued by another ad-network are being pooled by <num> potentially problematic publishers”.
Advertisers. Because we do not expect advertisers to be particularly technically adept, we frame our notification around the presence of their ad creatives on potentially problematic content as follows: “An ad creative associated with your brand was observed on <num> problematic publishers. This association could negatively impact your reputation and future business.”
Reports attached to notification emails. To provide concrete evidence to support the reported vulnerabilities, we generated an automated PDF report, with appropriate branding based on the notification source. This report contained: (1) up to 25 instances of vulnerabilities which involve the specific entity; (2) an explanation of the root causes and implications of each vulnerability; and (3) possible remediation options. In addition, notifications to advertisers also included attachments of screenshots showing the advertiser’s ad creative on a problematic publisher and the HAR file associated with the page load. Overall, similar notification framing was used for each stakeholder.
@@
Timeline of notifications. Figure 3 depicts the timeline of our notification campaign. Prior to each round of notifications we conduct static and dynamic analysis to identify dark pooling vulnerabilities w.r.t to our dataset of problematic publishers (cf. Section 3.1). Then, we generate and send notifications and reports for each entity associated with that round. If we do not hear back, we send a reminder email approximately 1-2 weeks after the initial email. We then wait a month before re-conducting our static and dynamic analysis. We use the data from this analysis to assess the influence of notifications sent during the current round.